OpenStudy (anonymous):

First person to answer correctly gets a medal What is the missing number in this sequence? 6 , -2 , -10 , -18, ___ , -34 A. -23 B. -26 C. -28 D. -25

OpenStudy (kagıtucak):

B

OpenStudy (solomonzelman):

the difference is -8 -18-2x=-34

OpenStudy (solomonzelman):

-18-8=-34+8=26

OpenStudy (solomonzelman):

i meant -26

OpenStudy (loser66):

free answer site?